OpenStudy (anonymous):

A stone is thrown in a vertically upward direction with a velocity of 5m s^-1. If acceleration of the stone in the downward direction 10m s^-2, what will be the height attained by the stone and how much time will it take to reach there?

OpenStudy (anonymous):

\[V_{int} = 5 ms^{-1}\] \[a_g = -10 ms^{-2} \] At the highest point \[V_{final} = 0\] \[V_{final} = V_{int} + a_gt\]

OpenStudy (anonymous):

or Use this one, this one is much better \[V_{final}^{2} - V_{int}^2 = 2a_g H\]

OpenStudy (anonymous):

sorry took wrong values \[ 25 = 20 \times H\] \[H = \frac{25}{20}\]

OpenStudy (anonymous):

this is the correct one

OpenStudy (anonymous):

did you get that ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think i got the correct answer is it s= 1.25 m and t= 0.5 sec?

OpenStudy (anonymous):

it is right ! hit the good answer

OpenStudy (anonymous):

First the time. \[V _{o}+a \times t =0\] So 5m/s-10m/s^2(t)=0 5m/s=10m/s^2(t) .5s=(t) \[D =(V _{o}+V)/2\times t\] D=(5m/s+0m/s)/2*.5=5m/s/4=1.25m

OpenStudy (anonymous):

ok..... u=5m/s a=10 m\(s^2) s=ut+(1/2)a(t^2) and t=usin(t)/g now t=90degree so t=5/10= 0.5s and s=5*0.5+(1/2)(10*(0.5^2)= 2.5+(0.5*0.25*10)=2.5+1.25= 3.75m

OpenStudy (anonymous):

ohh sorry since at height, u=0... so. s=(1/2)a(t^2)=(1/2)*10*(0.5^2)=1.25m